BROADBAND CHIRPED-PULSE FOURIER-TRANSFORM MICROWAVE SPECTROSCOPIC INVESTI-GATION OF THE STRUCTURES OF THREE DIETHYLSILANE CONFORMERS

\underline{AMANDA L. STEBER}, DANIEL A. OBENCHAIN, REBECCA A. PEEBLES, and SEAN A. PEE-BLES, Department of Chemistry, Eastern Illinois University, 600 Lincoln Avenue, Charleston, IL 61920; JUSTIN L. NEILL, MATT T. MUCKLE, and BROOKS H. PATE, Department of Chemistry, University of Virginia, Charlottesville, VA 22904; GAMIL A. GUIRGIS, Department of Chemistry and Biochemistry, The College of Charleston, Charleston, SC 29424.

The rotational spectrum of diethylsilane has been assigned using broadband chirped-pulse Fourier-transform microwave (CP-FTMW) spectroscopy. Previously, Fourier-transform microwave rotational spectra were observed using a Balle-Flygare type instrument for the ^{28}Si isotopologues of the gauche-gauche, trans-gauche, and trans-trans conformers.^{a} In the present study, a broadband microwave spectrum was obtained at the University of Virginia, taking advantage of the ability to perform deep signal averaging to increase the measurement sensitivity. To obtain a full structural determination of the conformers of this molecule, spectra for the ^{29}Si, ^{30}Si, and single ^{13}C substitutions for the gauche-gauche, the trans-gauche, and the trans-trans species were assigned. Substitution (r_{s}) structures and inertial fit (r_{0}) structures were determined and a comparison between the experimental and ab initio structures will be presented. For the ^{28}Si isotopologues, the percent differences between the experimental and ab initio rotational constants are less than 1.5% for the trans-trans and trans-gauche and are between 2.0 and 5.0% for the gauche-gauche conformer. The structural parameters will be compared between this molecule, diethylgermane and other silicon containing molecules and the relative abundances of the three conformers will be discussed.

aS.A. Peebles, M.M. Serafin, R.A. Peebles, G.A. Guirgis, and H.D. Stidham J. Phys. Chem. A, (2009), DOI: 10.1021/jp811049n.